a=(/3+/2)^-3b=(/3-/2)^-3Then(a+1)^-1 + (b+1)^-1 =?

Grade:12th pass

1 Answers

SOVON SARKAR
6 years ago
Let /3-/2=x. Therefore /3+/2=1/x
Now a=x^3 and b=1/x^3.  a+1=x^3+1........(1)
Now b+1=1/x^3 + 1=(x^3+1)/x^3=(a+1)/x^3   [using (1)]
Therefore 1/(b+1)=x^3/(a+1) 
Now 1/(a+1) + 1/(b+1) = 1/(a+1)+x^3/(a+1)=(x^3+1)/(a+1)=(a+1)/(a+1)=1  [using (1)]